# The 2010 Census Confidentiality Protections Failed, Here’s How and Why

### v1

@@ 8. Reidentification Results
Having demonstrated that it is possible to reconstruct an accurate image of HDF from the publicly released tables, we now analyze the results of experiments that characterize what an attacker could learn from these data. We want to assess the range of possible confidentiality breaches as a function of the quality of the attacker’s external information holding constant the quality of the microdata reconstruction from published tables and the record-linkage algorithms. The COMCRL database described in Section 3.5 has deficiencies in both coverage and accuracy relative to the CEF. Only $37.1\%$ of COMRCL records with valid {pik, block, sex, agebin} feature values matched an in-universe record on the CEF (Table 3). Because we use information from the CEF to characterize vulnerable populations and because confirmation requires that all values of {pik, block, sex, (age or agebin), race, ethnicity} match, this poor agreement between COMCRL and CEF limits the achievable putative and confirmed reidentification rates but does not affect reidentification precision rates. More accurate commercial databases exist today, and other forms of external data are also feasible attacker information—for example, government databases, web-scraped data, or informal knowledge from friendships or coworkers. To model an attacker with access to higher quality external data, we use the specially prepared extract from the CEF called $\text{CEF}_{atkr}$ described in Section 5 that includes only $\{pik,block,sex,age\}$ as the attacker’s data. [^28] While better informed than an attacker using COMCRL, the $\text{CEF}_{atkr}$ attacker is still not worst-case because we hold constant the methodology used to reconstruct HDF and match to the external data. In particular, we use a straightforward record-linkage attack algorithm without any enhancements. Using strong linkage models, including models that search for a match in nearby blocks, would certainly strengthen the attack. When using $\text{CEF}_{atkr}$ as the attacker’s data, all data-defined CEF records are potentially at risk of confirmation.
@@
To calibrate the reidentification rates from $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ as the source of {race, ethnicity}, we also analyze two types alternatives or baselines, each of which provide different insights. First, we use the CEF and HDF themselves as the source of {race, ethnicity}. Using these data as the source of the target characteristics demonstrates how accurate the attacker’s inferences would be if the confidential microdata themselves were released without any SDL treatment (CEF) or using only record-level SDL treatments (HDF) with a naive attacker that does not attempt to undo swaps. [^29] When the attacker’s external data is $\text{CEF}_{atkr}$, we only perform the experiment using the HDF as the source of {race, ethnicity}. Second, as described in Section 5.3, we also incorporate two statistical baselines as the source of {race, ethnicity}. The first, MDG, attaches the within-block modal values of {race, ethnicity} to all records in the block. The second, PRG, randomly assigns {race, ethnicity} in proportion to the share of each pair within the census block. Because the MDG and PRG baselines assign {race, ethnicity} from the HDF matching on {block, sex agebin} and because $\text{rHDF}_{b,t}$ perfectly replicates the {block, sex, agebin} feature values from HDF, the binned-age putative matches are identical for experiments using $\text{rHDF}_{b,t}$, $\text{rHDF}_{b}$, MDG, and PRG. [^30] When we use COMRCL records that match to the CEF on {pik, block}, putative reidentifications are not identical to those obtained from MDG and PRG.
@@
In Section 8.1 we present reideintification metrics for the data-defined national population by census block size. Despite demonstrating high precision on inferences about {race, ethnicity} from reidentifications resulting from augmenting an external database using the reconstructed HDF, it is difficult to draw conclusions about confidentiality breaches based on the national block-size statistics. Hence, we consider vulnerable populations in Section 8.2, which presents results for nonmodal persons and contrasts these results with those for modal persons. Then, in Section 8.3, we complete the discussion of vulnerable populations by considering nonmodal persons with 0 solution variability who are population uniques in the reconstructed HDF (and, hence in the HDF due to the 0 solution variability).
@@
While it is unclear which inferences about individuals are statistical versus privacy-breaching in the overall population, our analysis of vulnerable populations identifies inferences about {race, ethnicity} pairs that are clearly privacy breaches because the improvement in reidentification precision is due primarily (for nonmodal persons) or exclusively (for population unique, 0 solution variability, nonmodal persons) to the presence of the target individual’s data in the published tables. To illustrate the significance of this point, consider the data for the state of Montana in the 2010 SF1. According to SF1 Table P9, 87.8% of Montana residents in 2010 indicated Non-Hispanic White-alone as their race and ethnicity. From this fact alone we can make many high-quality inferences. For a randomly selected COMRCL record in Montana, we expect that record to be Non-Hispanic White-alone with about 88% probability. This does not, however, suggest that publishing a modest number of state-level statistics generates prohibited confidentiality violations because although these are inferences about individuals, they are precise because of racial and ethnic homogeneity in Montana and are, thus, predictable on a purely statistical basis. However, our results also show that the single person living in Montana who responded Not Hispanic or Latino, White, American Indian/Alaska Native, Asian, Native Hawaiian/Other Pacific Islander—a rare non-Hispanic 4-race response that is a 0 solution variability population unique on {block, sex, agebin, race, ethnicity}—can be reidentified with precision of 95.2%. That is, the “guess” that this person has the unusual non-Hispanic, 4-race feature value is wrong less than 1 time in 20. Had that person not responded to the 2010 Census, the precision would have been exactly 0 for the modal guesser and, approximately, $1/989,414$ or effectively 0.0% for the proportional guesser. That is, the attacker would “never have guessed” that response. The only reason this unusual person can be geolocated in a block of 15 persons and reidentified so precisely in the 2010 Census tables is that the actual response was used in all tables of SF1. Such inferences are confidentiality breaches.
@@
The replication archive contains an ExcelTM workbook (rhdf_bt_0solvar_extract.xlsx) with reconstructed records from 29 census tracts in which every block has 0 solution variability. Tags identify every row that is population unique on {block, sex, agebin} and {block, sex, agebin, race, ethnicity} in the spreadsheet. No access to confidential data is required to produce or confirm these tables, and any population unique on agebin must also be population unique on age. Every population unique in that spreadsheet is known with certainty to be a population unique in the confidential HDF. The only reason the reidentification precision is not 100% is the record-level SDL that was applied to CEF to create HDF. But that SDL treatment was not designed to protect (1) full enumeration microdata files, as opposed to a 10% microdata sample, (2) full enumeration microdata files containing geographic areas with populations as small as 1 person, as opposed to the 100,000 minimum required for microdata samples, or (3) full enumeration files containing {race, ethnicity} responses as rare as non-Hispanic, White, American Indian/Alaska Native, Asian, and Native Hawaiian/Other Pacific Islander, which has a national population of only 7,460 as opposed to the 10,000 national population minimum required for microdata samples [mckenna:2019:microdata]. This is a classic example of inconsistent SDL. Evaluated according to the 2010 Census tabular data rules [mckenna2018disclosure] the release is “safe.” Evaluated under the 2010 Census microdata rules, it fails catastrophically. But it is exactly the same data release expressed in two mathematically equivalent representations—34 tabular summaries shown in Table 2 and 308,745,538 microdata records using the same feature set shown in equation (2).
@@
Differentiating statistical or generalizable inferences from confidentiality-violating inferences can be subtle. As we described in Section 5.3, in the absence of computational or time constraints, an ideal experiment for estimating the distribution of empirical privacy loss in the 2010 Census publications would iteratively remove each individual’s record, treat it as a target, and repeat the entire set of reconstruction and reidentification experiments, examining inferences about that target person. The intuition behind this approach is especially clear in the case of reidentification. If the target person’s data were not in the confidential database then we could not assert that the feature values we attached to their record in the attacker’s data, either COMRCL or $\text{CEF}_{atkr}$, were the result of a prohibited reidentification. Indeed, as discussed in Section 2.1, the Census Act (13 U.S. Code §§ 8(b) & 9) specifically identifies the source of the confidential information protected by statute as “information reported by, or on behalf of, any particular respondent,” and “data furnished by any particular establishment or individual,” which reflects the legal focus on data present in internal response databases. The ideal experiment extends this concept, arguing that even if an inference is made about a target person whose record is in the CEF, this inference should not be regarded as confidentiality-violating if that same inference could have been made after removing the target person’s record from the CEF. But, as the Montana example makes clear, if that inference could only have been made because the target person’s record was present in the CEF, that is a confidentiality breach.
@@
Unfortunately, “leaving out” each record in the U.S. population, simulating the 2010 Census SDL treatments, then performing the reconstruction and reidentification attacks for each such record is prohibitively expensive in time and computational resources. To approximate the logic of this ideal experiment, however, we have focused on vulnerable populations—infrequently occurring records that would be very difficult or impossible to reidentify if not for their participation in the 2010 Census. Specifically, we focus on records with {race, ethnicity} not equal to the modal value within their block, and which are unique in {block, sex, agebin}. We also further subset to cases for which the attacker can be confident that the reconstructed microdata exactly match the confidential HDF—blocks with 0 solution variability. Accurate inferences about such records in our our experiments would be much more difficult, and often impossible, had the person’s record not been present in the CEF.
@@ 8.1. National Results
Table 9 presents the statistics for putative and confirmed reidentifications, and for precision rates for all data-defined persons in the 2010 Census for all data-defined persons in COMRCL and $\text{CEF}_{atkr}$. Detailed results by census block size appear in Appendix Table 12, the top panel of which is identical to Table 9. The results in Table 9 are refinements of those first released in 2019 [abowd:2019:AAAS]. They show that when the attacker uses the low quality COMRCL data, the reconstruction $\text{rHDF}_{b,t}$ produces 166,100,000 putative reidentifications (57.9% putative rate for the data-defined population) of which 68,480,000 are confirmed (23.9% confirmation rate for the data-defined population) with a reidentification precision rate of 41.2%. When the attacker uses the high-quality quasi-identifier data in $\text{CEF}_{atkr}$, the reconstruction $\text{rHDF}_{b,t}$ yields 267,800,000 putative reidentifications (97.0% of the data-defined population) of which 208,500,000 are confirmed (75.5% of the data-defined population) with a precision rate of 77.9%.
@@
Alg. 3) Size (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%)
@@
significant digits to conform to disclosure limitation requirements. Population for attacker COMRCL
@@
is the total number of data-defined records in COMRCL that are also in the CEF universe (see
@@
Table 3). Population for attacker $\text{CEF}_{atkr}$ is the total number of data-defined CEF records.
@@
The Census Bureau released similar reidentification results as part of litigation surrounding the 2020 Census and to its Scientific Advisory Committee between 2019 and 2022 [abowd2_fairlines_appendixB, hawes:2022, abowd:hawes:2023] based on earlier versions of the models presented in this paper. Some researchers noted that baselines similar to MDG and PRG could produce reidentifications apparently comparable to those of $\text{rHDF}_{b,t}$ [ruggles:vanriper:2022, francis:2022]. The rows MDG and PRG in the COMRCL panel confirm that claim but only, as we shall see, if one ignores nonmodal and vulnerable populations. By construction, these two strategies have the same overall putative reidentifications as $\text{rHDF}_{b,t}$, and their confirmation and precision rates are also comparable. Comparison to statistical baselines at the national level highlights the fact that national reidentification rates, while informative about the scale of the match (e.g., 68,480,000 individuals had confirmed {race, ethnicity} in the COMRCL experiment using $\text{rHDF}_{b,t}$), are difficult to interpret in terms of privacy loss. The misleading similarity between statistical baselines and $\text{rHDF}_{b,t}$ requires using leave-one-out reasoning to resolve, which we do below. $\text{rHDF}_{b,t}$ is able to reidentify the race and ethnicity of nonmodal individuals that are unique in the linking quasi-identifiers {block, sex, agebin} in blocks with 0 solution variability, where MDG and PRG completely fail—a much clearer demonstration of privacy loss. [^31] Since this paper is primarily a proof of concept, a larger-scale attack using more tables and reconstructing additional features would result in less solution variability and more population uniques.
@@
Surprisingly, releasing the CEF itself with the full feature set in Table 1 (except pik) would produce only relatively modest increase in the putative reidentification rates compared to HDF, $\text{rHDF}_{b,t}$, $\text{rHDF}_{b}$ or either baseline, as shown in the first row of Table 9 for both COMCRL and $\text{CEF}_{atkr}$. As noted in Table 3, COMCRL data did not align well with the data collected in the 2010 Census. But even if the attacker used high-quality input data, as in $\text{CEF}_{atkr}$, releasing the CEF (except pik) with the feature set in Table 1 would still produce only modestly more putative reidentifications, and releasing the HDF with the feature sets shown in Table 1 would produce exactly the same putative reidentifications, as $\text{rHDF}_{b,t}$, $\text{rHDF}_{b}$, or either baseline. From the viewpoint of records at risk for record-linkage reidentification, the reconstructed microdata and the statistical baselines accurately replicate HDF. The confirmation and precision rates for CEF and HDF are marginally higher than those for $\text{rHDF}_{b,t}$, $\text{rHDF}_{b}$ and MDG, and substantially better than PRG.

### v2

@@ 7. Reidentification Results
Having demonstrated that it is possible to reconstruct an accurate image of HDF from the publicly released tables, we now analyze the results of experiments that characterize what an attacker could learn from these data. We want to assess the range of possible confidentiality breaches as a function of the quality of the attacker’s external information holding constant the quality of the microdata reconstruction from published tables and the record-linkage algorithms. The COMRCL database (see Section 3.5) has deficiencies in both coverage and accuracy relative to the CEF. As indicated in Table 3, only $37.1\%$ of COMRCL records with valid {pik, block, sex, agebin} feature values matched an in-universe record on the CEF. Because we use information from the CEF to characterize vulnerable populations and because confirmation requires that all values of {pik, block, sex, (age or agebin), race, ethnicity} match, this poor agreement between COMRCL and CEF limits the achievable putative and confirmed reidentification rates but does not affect reidentification precision rates. More accurate commercial databases exist today, and other forms of external data are also feasible attacker information—for example, government databases, web-scraped data, or informal knowledge from friendships or coworkers. To model an attacker with access to higher quality external data, we use the specially prepared extract from the CEF called $\text{CEF}_{atkr}$ that includes only $\{pik,block,sex,age\}$ as the attacker’s data (see Section 5). [^30] While better informed than an attacker using COMRCL, the $\text{CEF}_{atkr}$ attacker is still not worst-case because we hold constant the methodology used to reconstruct HDF and match to the external data. In particular, we use a straightforward record-linkage attack algorithm without any enhancements. Using strong linkage models, including models that search for a match in nearby blocks, would certainly strengthen the attack. When using $\text{CEF}_{atkr}$ as the attacker’s data, all data-defined CEF records are potentially at risk of confirmation.
@@
To calibrate the reidentification rates from $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ as the source of {race, ethnicity}, we also analyze two types of alternatives or baselines, each of which provide different insights. First, we use the CEF and HDF themselves as the source of {race, ethnicity}. Using these data as the source of the target characteristics demonstrates how accurate the attacker’s inferences would be if the confidential microdata themselves were released without any SDL treatment (CEF) or using only record-level SDL treatments (HDF) with a naive attacker that does not attempt to undo swaps. [^31] When the attacker’s external data is $\text{CEF}_{atkr}$, we only perform the experiment using the HDF as the source of {race, ethnicity}. Second, we also incorporate two statistical baselines as the source of {race, ethnicity} (see Section 5.3). The first, MDG, attaches the within-block modal values of {race, ethnicity} to all records in the block. The second, PRG, randomly assigns {race, ethnicity} in proportion to the share of each pair within the census block. Because the MDG and PRG baselines assign {race, ethnicity} from the HDF matching on {block, sex agebin} and because $\text{rHDF}_{b,t}$ perfectly replicates the {block, sex, agebin} feature values from HDF, the binned-age putative matches are identical for experiments using $\text{rHDF}_{b,t}$, $\text{rHDF}_{b}$, MDG, and PRG. [^32] When we use COMRCL records that match to the CEF on {pik, block}, putative reidentifications are not identical to those obtained from MDG and PRG.
@@
We next present reidentification metrics for the data-defined national population by census block size (see Section 7.1). Despite demonstrating high precision on inferences about {race, ethnicity} from reidentifications resulting from augmenting an external database using the reconstructed HDF, it is difficult to draw conclusions about confidentiality breaches based on the national block-size statistics. Hence, we consider vulnerable populations by presenting results for nonmodal persons and contrasting these results with those for modal persons (see Section 7.2). We complete the discussion of vulnerable populations by considering nonmodal persons with zero solution variability who are population uniques in the reconstructed HDF and, hence, also unique in the HDF due to the zero solution variability (see Section 7.3). While it is unclear which inferences about individuals are statistical versus confidentiality-breaching in the overall population, our analysis of vulnerable populations identifies inferences about {race, ethnicity} pairs that are clearly confidentiality breaches because the improvement in reidentification precision is due primarily (for nonmodal persons) or exclusively (for population unique, zero solution variability, nonmodal persons) to the presence of the target individual’s data in the published tables.
@@
To illustrate the significance of this point, consider the data for the state of Montana in the 2010 SF1. According to SF1 Table P9, 87.8% of Montana residents in 2010 indicated non-Hispanic White alone as their race and ethnicity. From this fact alone we can make many high-quality inferences. For a randomly selected COMRCL record in Montana, we expect that record to be non-Hispanic White alone with about 88% probability. This does not, however, suggest that publishing a modest number of state-level statistics generates prohibited confidentiality violations because although these are inferences about individuals, they are precise because of racial and ethnic homogeneity in Montana and are, thus, predictable on a purely statistical basis. However, our results also show that the single person living in Montana who responded Not Hispanic or Latino, White, American Indian/Alaska Native, Asian, Native Hawaiian/Other Pacific Islander—a rare non-Hispanic 4-race response that is a zero solution variability population unique on {block, sex, agebin, race, ethnicity}—can be reidentified with precision of 95.2%. That is, the “guess” that this person has the unusual non-Hispanic, 4-race feature value is wrong less than 1 time in 20. Had that person not responded to the 2010 Census, the precision would have been exactly zero for the modal guesser and, approximately, $1/989,414$ or effectively 0.0% for the proportional guesser. That is, the attacker would “never have guessed” that response. The only reason this unusual person can be geolocated in a block of 15 persons and reidentified so precisely in the 2010 Census tables is that the actual response was used in all tables of SF1. Such inferences are confidentiality breaches.
@@
The replication archive contains an ExcelTM workbook (rhdf_bt_0solvar_extract.xlsx) with reconstructed records from 29 census tracts in which every block has zero solution variability. Tags identify every row that is population unique on {block, sex, agebin} and {block, sex, agebin, race, ethnicity} in the spreadsheet. No access to confidential data is required to produce or confirm these tables, and any population unique on agebin must also be population unique on age. Every population unique in that spreadsheet is known with certainty to be a population unique in the confidential HDF. The only reason the reidentification precision is not 100% is the record-level SDL that was applied to CEF to create HDF. But that SDL treatment was not designed to protect (1) full enumeration microdata files, as opposed to a 10% microdata sample, (2) full enumeration microdata files containing geographic areas with populations as small as 1 person, as opposed to the 100,000 minimum required for microdata samples, or (3) full enumeration files containing {race, ethnicity} responses as rare as non-Hispanic, White, American Indian/Alaska Native, Asian, and Native Hawaiian/Other Pacific Islander, which has a national population of only 7,460 as opposed to the 10,000 national population minimum required for microdata samples (mckenna:2019:microdata). This is a classic example of inconsistent SDL. Evaluated according to the 2010 Census tabular data rules (mckenna:2018:disclosure) the release is “safe.” Evaluated under the 2010 Census microdata rules, it fails catastrophically. But it is exactly the same data release expressed in two mathematically equivalent representations—34 tabular summaries shown in Table 2 and 308,745,538 microdata records using the same feature set shown in Equation 2.
@@
Differentiating statistical or generalizable inferences from confidentiality-violating inferences can be subtle. In the absence of computational or time constraints, an ideal experiment for estimating the distribution of empirical privacy loss in the 2010 Census publications would iteratively remove each individual’s record, treat it as a target, and repeat the entire set of reconstruction and reidentification experiments, examining inferences about that target person (see Section 5.3). The intuition behind this approach is especially clear in the case of reidentification. If the target person’s data were not in the confidential database then we could not assert that the feature values we attached to their record in the attacker’s data, either COMRCL or $\text{CEF}_{atkr}$, were the result of a prohibited reidentification. Indeed, the Census Act (13 U.S. Code §§ 8(b) & 9) specifically identifies the source of the confidential information protected by statute as “information reported by, or on behalf of, any particular respondent,” and “data furnished by any particular establishment or individual,” which reflects the legal focus on data present in internal response databases (see Section 2.2). The ideal experiment extends this concept, arguing that even if an inference is made about a target person whose record is in the CEF, this inference should not be regarded as confidentiality-violating if that same inference could have been made after removing the target person’s record from the CEF. But, as the Montana example makes clear, if that inference could only have been made because the target person’s record was present in the CEF, that is a confidentiality breach.
@@
Unfortunately, “leaving out” each record in the U.S. population, simulating the 2010 Census SDL treatments, then performing the reconstruction and reidentification attacks for each such record is prohibitively expensive in time and computational resources. To approximate the logic of this ideal experiment, however, we have focused on vulnerable populations—infrequently occurring records that would be very difficult or impossible to reidentify if not for their participation in the 2010 Census. Specifically, we focus on records with {race, ethnicity} not equal to the modal value within their block, and which are unique in {block, sex, agebin}. We also further subset to cases for which the attacker can be confident that the reconstructed microdata exactly match the confidential HDF—blocks with zero solution variability. Accurate inferences about such records in our our experiments would be much more difficult, and often impossible, had the person’s record not been present in the CEF.
@@ 7.1. National results
Table LABEL:tab:reid_national presents the statistics for putative reidentifications (numerator in Equation 11), confirmed reidentifications (numerator in Equation 12), and precision rates (Equation 13) for all data-defined persons in the 2010 Census for all data-defined persons in COMRCL and $\text{CEF}_{atkr}$. Detailed results by census block size appear in Appendix Table LABEL:tab:reid_blockpop, the top panel of which is identical to Table LABEL:tab:reid_national. The results in Table LABEL:tab:reid_national are refinements of those first released in 2019 (abowd:2019:AAAS). They show that when the attacker uses the lower-quality COMRCL data, the reconstruction $\text{rHDF}_{b,t}$ produces 166,100,000 putative reidentifications (57.9% putative rate for the data-defined population) of which 68,480,000 are confirmed (23.9% confirmation rate for the data-defined population) with a reidentification precision rate of 41.2%. When the attacker uses the high-quality quasi-identifier data in $\text{CEF}_{atkr}$, the reconstruction $\text{rHDF}_{b,t}$ yields 267,800,000 putative reidentifications (97.0% of the data-defined population) of which 208,500,000 are confirmed (75.5% of the data-defined population) with a precision rate of 77.9%.
@@
Alg. 3) Size (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%) (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%)
@@
significant digits to conform to disclosure limitation requirements. Population for attacker
@@
COMRCL is the total number of data-defined records in COMRCL that are also in the CEF
@@
universe (see Table 3). Population for attacker $\text{CEF}_{atkr}$ is the total number of data-
defined CEF records.
@@
The Census Bureau released similar reidentification results as part of litigation surrounding the 2020 Census and to its Scientific Advisory Committee between 2019 and 2022 (abowd2_fairlines_appendixB; hawes:2022; abowd:hawes:2023) based on earlier versions of the models presented in this paper. Some researchers noted that baselines similar to MDG and PRG could produce reidentifications apparently comparable to those of $\text{rHDF}_{b,t}$ (ruggles:vanriper:2022; francis:2022). The rows MDG and PRG in the COMRCL panel confirm that claim but only, as we shall see, if one ignores nonmodal and vulnerable populations. By construction, these two strategies have the same overall putative reidentifications as $\text{rHDF}_{b,t}$, and their confirmation and precision rates are also comparable. Comparison to statistical baselines at the national level highlights the fact that national reidentification rates, while informative about the scale of the match (e.g., 68,480,000 individuals had confirmed {race, ethnicity} in the COMRCL experiment using $\text{rHDF}_{b,t}$), are difficult to interpret in terms of privacy loss. The misleading similarity between statistical baselines and $\text{rHDF}_{b,t}$ requires using leave-one-out reasoning to resolve, which we do below. $\text{rHDF}_{b,t}$ is able to reidentify the race and ethnicity of nonmodal individuals that are unique in the linking quasi-identifiers {block, sex, agebin} in blocks with zero solution variability, where MDG and PRG completely fail—a much clearer demonstration of privacy loss. [^33] Since this paper is primarily a proof of concept, a larger-scale attack using more tables and reconstructing additional features would result in less solution variability and more population uniques.
@@
Surprisingly, releasing the CEF itself with the full feature set in Table 1 (except pik) would produce only relatively modest increase in the putative reidentification rates compared to HDF, $\text{rHDF}_{b,t}$, $\text{rHDF}_{b}$ or either baseline, as shown in the first row of Table LABEL:tab:reid_national for both COMRCL and $\text{CEF}_{atkr}$. As noted in Table 3, COMRCL data did not align well with the data collected in the 2010 Census. But even if the attacker used high-quality input data, as in $\text{CEF}_{atkr}$, releasing the CEF (except pik) with the feature set in Table 1 would still produce only modestly more putative reidentifications, and releasing the HDF with the feature sets shown in Table 1 would produce exactly the same putative reidentifications, as $\text{rHDF}_{b,t}$, $\text{rHDF}_{b}$, or either baseline. From the viewpoint of records at risk for record-linkage reidentification, the reconstructed microdata and the statistical baselines accurately replicate HDF. The confirmation and precision rates for CEF and HDF are marginally higher than those for $\text{rHDF}_{b,t}$, $\text{rHDF}_{b}$ and MDG, and substantially better than PRG.